How many gallons of cream that is 22% butterfat must be mixed with milk that is 2% buttermilk to get 20 gallons of milk containing 4% buttermilk?
Answer by ankor@dixie-net.com(22740)

How many gallons of cream that is 22% butterfat must be mixed with milk that is 2% buttermilk to get 20 gallons of milk containing 4% buttermilk?
:
Let x = the amt of 22% cream required
:
the resulting amt required is 20 gal, therefore
2% amt = (20-x)
:
.22x + .02(20-x) = .04(20); a simple equation
:
.22x + .40 - .02x = .80; multiplied what's inside the brackets
:
.22x - .02x = .80 - .40
:
.20 = .40
:
x = .40/.20
:
x = 2 gal of 22% cream required
Then
20 - 2 = 18 gal of 2% milk required
:
:
Check solution:
.22(2) + .02(18) = .04(20)
.44 + .36 =.80; confirms out solution